OpenStudy (anonymous):

y'' + y = 0 Can anyone explain to me how the solution to this ODE is c1sin(x)+c2cos(x)?

myininaya (myininaya):

r^2+1=0 r=i or r=-i

OpenStudy (anonymous):

Ohh okay and then thats put into Eulers simplification for complex roots? I see, I was taking that characteristic equation as r^2 + r, my mistake Thank you

myininaya (myininaya):

oh yes r^2+r would screw it up here y''+y'=0 you could do r^2+r=0

OpenStudy (anonymous):

Haha kk, thanks